# Quantifying Risks in Multi-turn Conversation with Large Language Models

### v1

@@ 4.3 Certification Results
We evaluate the certification bounds of several state-of-the-art large language models: Llama-3.3-70b-Instruct (Meta AI, 2024), Mistral-Large-2407 (Mistral AI, 2024), DeepSeek-R1 (Guo et al., 2025), gpt-oss-120b (Agarwal et al., 2025), and Claude-Sonnet-4 (Anthropic, 2024). We use the default hyperparameter settings shown in Table 3 and analyze their influence through an ablation study (Appendix C). For each LLM and specification, we estimate certification bounds on the attack success probability with $95\%$ confidence, reporting the median of the lower and upper bounds across all specifications under a distribution in Table 1. Figure 4 and 5 (Appendix B) show the results in box plots for specifications developed from the chemical_biological and cybercrime datasets respectively.
Table 1: Certification bounds under different distributions for each dataset and model (median of 95% confidence intervals across all specifications under a distribution). Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among four distributions for each LLM.
@@
By comparing the bounds, we observe that among frontier LLMs, Claude-Sonnet-4 is consistently safer than the others, while Mistral-Large and DeepSeek-R1 exhibit higher risks. In particular, DeepSeek-R1 reaches a certified lower bound of over 70% in cybercrime scenarios under RNwJ distributions. For LLMs with relatively low probabilities of catastrophic outcomes (e.g., Claude-Sonnet-4 and gpt-oss-120b), distributions augmenting with jailbreak are largely ineffective. In contrast, weaker LLMs such as Mistral-Large and DeepSeek-R1 remain vulnerable to jailbreak prompts, indicating that additional safety training is needed. We analyze the effect of the jailbreak probability in Appendix C.5; for less safe LLMs, increasing the jailbreak probability generally raises catastrophic outcomes, while for safer LLMs the effect is negligible. Other distributions, Adaptive with Rejection and Graph Path, are often more effective in producing catastrophic outcomes on safer LLMs. For Graph Path, constraining the final query to a harmful set (GPh) consistently increases attack effectiveness relative to the vanilla last-query distribution (GPv), which shows that shaping the final step of a multi-query sequence is an effective method for attackers. For Adaptive with Rejection, the strategy exploits the fact that safer LLMs refuse to answer queries at non-trivial rates (roughly 20% for gpt-oss-120b and 15% for Claude-Sonnet-4 in our samples). By designing sequences that interact with these rejection dynamics, attackers can substantially increase catastrophic responses on LLMs that otherwise appear well aligned.
Attack Patterns.
In our analysis, we identify two common attack patterns:
  1. 1.
    Effect of Distractors. We observe that LLMs often refuse to answer harmful queries that are presented in isolation, replying with messages such as “I can’t provide that information” due to built-in safety mechanisms (Zhang et al., 2025; Yuan et al., 2025). However, when the same query is embedded in a multi-turn dialogue that includes other questions, even not directly related, the model is more likely to provide a harmful answer. This behavior is observed across our various specification distributions. Figure 2 shows a representative example of specification in the graph path (harmful target constraint) distribution.
  2. 2.
Comparing with Baselines
Table 2: Numbers of attack scenarios where the certification lower bound exceeds the baseline ST(Single-turn) and MT(Multi-turn) ASR by more than 0.05.
@@
The baselines are not directly comparable, but in the absence of stronger alternatives, we provide a rough comparison. Importantly, our certification evaluates models over a much larger conversation space, considering all possible sequences consistent with the query distributions rather than a fixed subset. To make the comparison more meaningful, for these baselines, in each scenario, we measure the fraction of queries (ST) or sequences (MT) that lead to catastrophic responses. Rather than using a binary outcome per scenario (recording a $1$ if any catastrophic response occurs across several trials, which is commonly done in the literature (Zou et al., 2023; Qi et al., 2023)), this measure provides a finer-grained view of how difficult it is to elicit catastrophic outcomes from a model in a given scenario. We then count the number of scenarios where our certified lower bound exceeds this baseline fraction by more than $0.05$ (Table 2).
@@ 5 Conclusion
Our work is the first to provide quantitative certificates for catastrophic risks in multi-turn LLM conversations. It can help model developers systematically evaluate and compare their models before deployment, and inform the general public of potential harms when interacting with LLMs. Since QRLLM only requires black-box access, it applies equally to both open- and closed-source models, thus broadening its utility.
@@ Appendix A Explicit Jailbreak Distribution
This defines $\mathcal{D}_{{jb}}$, which is used in the main analysis as a black-box sampling distribution.
Appendix B Detailed Certification Bounds
Figure 4 and  5 report the complete certification lower and upper bounds (median and IQR) for every model–distribution pair across all specifications.
@@ Appendix B Detailed Certification Bounds
Appendix C ablation study
In this section, we analyze the effect of hyperparameters on certification results. Table  3 shows the hyperparameters and their values used in the experiments. We conduct ablation studies on a randomly selected scenario from the dataset on Graph Path (harmful target constraint) distribution. Unless otherwise stated, we certify Llama-3.3-70b-instruct.
@@
C.1 Number of Samples
To assess how our certification bounds change with the number of samples $n$, we report bounds in Figure 8(a). The ranges between lower and upper bounds shrink as $n$ increases from small values, and stablize once $n\approx 50$. In our main experiments, we therefore adopt $n=50$ as a trade off between computational cost and statistical precision.
C.2 Ratio of Weight
To study the influence of this ratio, we perform an ablation experiment by varying $\lambda_{h}/\lambda_{l}$ across the values $\{1.5,2.0,2.5,3.0,3.5\}$. Note that we require $\lambda_{h}>\lambda_{l}$, hence the minimum ratio considered is $1.5$. We then evaluate the resulting certified bounds under these different settings. As shown in Figure 8(b), both the lower and upper bounds generally increase as the ratio grows, suggesting that placing relatively more weight on the high-weight neighbor set improves the effectiveness of this distribution.
C.3 Graph Thresholds
((a))
((b))
Figure 6: Ablation studies for certification bounds under different graph-threshold settings.
C.4 Length of Query Sequence
Figure 8(c) shows how certification bounds vary with the length of query sequence. Increasing the length consistently shifts the bounds upward, suggesting that longer query sequences have higher probability of trigger catastrophic outcomes on models.
C.5 Jailbreak Probability
Certification bounds on Random Node with Jailbreak distribution is controlled by the jailbreak probability hyperparameter. We study three LLMs: Llama-3.3-70B-Instruct, Deepseek-R1, and Claude-Sonnet-4 and show results in Figure 7. We find that the effect of jailbreak probability is model-dependent: for Llama-3.3-70B-instruct, bounds decrease slightly when the probability is $1$, but increase at lower values, suggesting that overly frequent jailbreak attempts can be detected. For Deepseek-R1, which is less robust, bounds grow steadily as the probability increases. For Claude-Sonnet-4, the bounds remain largely unchanged with values close to 0, indicating that it is defensive to jailbreak attempts.
((a))
((b))
((c))
Figure 7: Ablation studies for jailbreak probability on certification bounds for three LLMs.
C.6 Size of Query Set
We show the effect of the size of the query set we build specifications on. Figure 8(d) shows that increasing the set size, i.e., the queries contained in the set, has little impact on the bounds, indicating that our results can generalize well to larger datasets.
@@ C.6 Size of Query Set
Figure 8: Ablation studies for (a) number of samples, (b) ratio of weights, (c)length of query sequence, (d) size of the query set, (e) variance.
C.7 Variance
We show the variance of our certification bounds in Figure 8(e), where we run the same experiment on one specification 10 times. We report the median and interquartile range (IQR) of the resulting 95% confidence lower and upper bounds. The results demonstrate that the variance is low, demonstrating the reliability of our certification procedure.
Appendix D LLM Usage

### v2

@@ 4.3 Certification Results
We evaluate the statistical certification bounds of several state-of-the-art large language models: Llama-3.3-70b-Instruct (Meta AI, 2024), Mistral-Large-2407 (Mistral AI, 2024), DeepSeek-R1 (Guo et al., 2025), gpt-oss-120b (Agarwal et al., 2025), Claude-Sonnet-4 (Anthropic, 2024) and Nova Premier (Intelligence, 2025). We use the default hyperparameter settings shown in Table 6 and analyze their influence through an ablation study (Appendix D). For each LLM and specification, we estimate statistical certification bounds on the attack success probability with $95\%$ confidence, reporting the median of the lower and upper bounds on chemical_biological and cybercrime categories across all specifications under a distribution in Table 1. Figure 4 and 5 (Appendix B) show the results in box plots for specifications developed from the chemical_biological and cybercrime datasets respectively.
Table 1: Statistical certification bounds under different distributions for each dataset and model (median of 95% confidence intervals across all specifications under a distribution). Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among four distributions for each LLM.
@@
By comparing the bounds, we observe that among frontier LLMs, Claude-Sonnet-4 and Nova Premier are safer than the others, while Mistral-Large and DeepSeek-R1 exhibit higher risks. In particular, Nova Premier demonstrates consistently low risk levels, largely because its built-in guardrails often block potentially unsafe content. On the other hand, DeepSeek-R1 reaches a certified lower bound of over 70% in cybercrime scenarios under RNwJ distributions. For LLMs with relatively low probabilities of catastrophic outcomes (e.g., Nova Premier and Claude-Sonnet-4), distributions augmenting with jailbreak are largely ineffective. In contrast, weaker LLMs such as Mistral-Large and DeepSeek-R1 remain vulnerable to jailbreak prompts, indicating that additional safety training is needed. We analyze the effect of the jailbreak probability in Appendix D.1; for less safe LLMs, increasing the jailbreak probability generally raises catastrophic outcomes, while for safer LLMs the effect is negligible. Other distributions, Adaptive with Rejection and Graph Path, are often more effective in producing catastrophic outcomes on safer LLMs. For Graph Path, constraining the final query to a harmful set (GPh) consistently increases attack effectiveness relative to the vanilla last-query distribution (GPv), which shows that shaping the final step of a multi-query sequence is an effective method for attackers. For Adaptive with Rejection, the strategy exploits the fact that safer LLMs refuse to answer queries at non-trivial rates (roughly 20% for gpt-oss-120b and 15% for Claude-Sonnet-4 in our samples). By designing sequences that interact with these rejection dynamics, attackers can substantially increase catastrophic responses on LLMs that otherwise appear well aligned.
Table 2: Numbers of attack scenarios where the statistical certification lower bound exceeds the baseline ST(Single-turn) and MT(Multi-turn) ASR by more than 0.05.
@@
nova 44 0
Attack Patterns.
In our analysis, we identify two common attack patterns:
  1. 1.
    Effect of Distractors. We observe that LLMs often refuse to answer harmful queries that are presented in isolation, replying with messages such as “I can’t provide that information” due to built-in safety mechanisms (Zhang et al., 2025; Yuan et al., 2025).
    However, when the same query is embedded in a multi-turn dialogue that includes other questions, even not directly related, the model is more likely to provide a harmful answer. This behavior is observed across our various specification distributions. Figure 2 shows a representative example of specification in the graph path (harmful target constraint) distribution.
  2. 2.
Comparing with Baselines
@@
The baselines are not directly comparable, but in the absence of stronger alternatives, we provide a rough comparison. Importantly, our statistical certification evaluates models over a much larger conversation space, considering all possible sequences consistent with the query distributions rather than a fixed subset. To make the comparison more meaningful, for these baselines, in each scenario, we measure the fraction of queries (ST) or sequences (MT) that lead to catastrophic responses. Rather than using a binary outcome per scenario (recording a $1$ if any catastrophic response occurs across several trials, which is commonly done in the literature (Zou et al., 2023; Qi et al., 2023)), this measure provides a finer-grained view of how difficult it is to elicit catastrophic outcomes from a model in a given scenario. We then count the number of scenarios where our certified lower bound exceeds this baseline fraction by more than $0.05$ (Table 2).
@@ 5 Conclusion
Our work is the first to provide quantitative certificates for catastrophic risks in multi-turn LLM conversations. It can help model developers systematically evaluate and compare their models before deployment, and inform the general public of potential harms when interacting with LLMs. Since C3LLM only requires black-box access, it applies equally to both open- and closed-source models, thus broadening its utility.
@@
Acknowledgements
This work was supported by a research gift from Amazon Research.
@@ Appendix A Explicit Jailbreak Distribution
This defines a base distribution over jailbreak-prefix strings, which we denote by $\mathcal{D}_{\mathrm{prefix}}$.
Augmentation with mutations and insertion. Let $\mathcal{D}_{\mathrm{prefix}}$ be the base distribution over jailbreak-prefix strings $g$, and let $\mathrm{tok}(g)=(t_{1},\dots,t_{L})$ be the tokenization of $g$. For a fixed mutation probability $\mu\in(0,1)$ and a replacement distribution $q$ over the tokenizer vocabulary (e.g., uniform or restricted to a set of “possible” tokens), we define the mutation operator $M_{\mu}$ by
$\Pr(\tilde{g}\mid g)\;=\;\prod_{i=1}^{L}\Bigl[(1-\mu)\,\mathbf{1}\{\tilde{t}_{i}=t_{i}\}\;+\;\mu\,q(\tilde{t}_{i})\Bigr],$
[…]
$\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}}(\tilde{g})\;=\;\sum_{g}\mathcal{D}_{\mathrm{prefix}}(g)\,\Pr(\tilde{g}\mid g).$
Given a base query $v$ and a jailbreak insertion probability $p\in(0,1)$, we define the jailbreak augmentation distributions over full queries by
$\mathcal{D}_{\mathrm{jb}}(a\mid v)\;=\;(1-p)\,\mathbf{1}\{a=v\}\;+\;p\sum_{g}\mathcal{D}_{\mathrm{prefix}}(g)\,\mathbf{1}\{a=g\mathbin{\|}v\},$
and
$\mathcal{D}_{\mathrm{jb}}^{\mathrm{mut}}(a\mid v)\;=\;(1-p)\,\mathbf{1}\{a=v\}\;+\;p\sum_{\tilde{g}}\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}}(\tilde{g})\,\mathbf{1}\{a=\tilde{g}\mathbin{\|}v\},$
where $g\mathbin{\|}v$ denotes the concatenation of the prefix $g$ and the base query $v$. Equivalently, $\mathcal{D}_{\mathrm{jb}}(\cdot\mid v)$ and $\mathcal{D}_{\mathrm{jb}}^{\mathrm{mut}}(\cdot\mid v)$ can be implemented by sampling a Bernoulli random variable $B\sim\mathrm{Bernoulli}(p)$ and setting
$a=\begin{cases}v,&B=0,\\[2.84526pt] g\mathbin{\|}v,&B=1,\;\;g\sim\mathcal{D}_{\mathrm{prefix}},\\ \end{cases}\quad\text{or}\quad a=\begin{cases}v,&B=0,\\[2.84526pt] \tilde{g}\mathbin{\|}v,&B=1,\;\;\tilde{g}\sim\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}},\end{cases}$
respectively.
Illustrative Scenario In our added experiments, we instantiate a context-dependent augmentation for the Adaptive with Rejection (AwR) attack: whenever the previous query is rejected, we replace the next base query $v_{t}$ with an augmented query $\tilde{v}_{t}\sim\mathcal{D}^{\mathrm{mut}}_{\mathrm{jb}}(\cdot\mid v_{t})$ (Appendix A). Table 3 reports certified 95% confidence intervals for catastrophic failure probability on one representative scenario.
Table 3: Certified 95% confidence intervals for catastrophic failure probability on a single representative scenario under the augmented adaptive-with-rejection (AwR) attack.
Model 95% CI
Mistral (0.7569, 0.9547)
gpt-oss (0.1463, 0.4034)
Claude (0.2121, 0.4877)
DeepSeek (0.4518, 0.7359)
Llama (0.4518, 0.7359)
Appendix B Detailed Statistical Certification Bounds
Figure 4 and  5 report the complete statistical certification lower and upper bounds (median and IQR) for every model–distribution pair across all specifications.
@@ Appendix B Detailed Statistical Certification Bounds
Appendix C Additional Results on HarmBench illegal category
Setup.
In addition to the chemical_biological and cybercrime categories reported in the main paper, we evaluate our certification framework on scenarios from the HarmBench (Mazeika et al., 2024) illegal subset. We follow the same graph construction procedure (Section 4.1) and the same experiment setup (Section 4.2). Due to time and compute constraints, we evaluate 30 scenarios for Claude, gpt-oss, and Llama, and the first 10 scenarios for Mistral and DeepSeek. We report results separately in Tables 4 and 5.
Distributions.
We evaluate the same four distributions used throughout the paper: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). Following the main text, each entry in Tables 4 and 5 is the median of 95% confidence intervals across all specifications under a distribution, and we bold the highest bound among the four distributions for each model.
Results and observations.
Compared with the main results, certified bounds on the illegal category for gpt-oss and Claude are lower than on the chemical_biological and cybercrime categories. By contrast, Llama, Mistral and DeepSeek exhibit high certified catastrophic risk across all three domains. These results illustrate how our framework allows practitioners to assess catastrophic risk with statistical guarantees in different domains and to choose models accordingly: for instance, Claude may be acceptable for illegal in our setting, whereas for cybercrime none of the evaluated models appears reliably safe, indicating that stronger mitigations on safer LLMs would be needed.
Table 4: Statistical certification bounds on HarmBench illegal (first 10 scenarios), Mistral and DeepSeek. Entries are the median of 95% confidence intervals across all specifications under a distribution. Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among the four distributions for each model.
Model RNwJ GPv GPh AwR
mistral (0.586, 0.845) (0.234, 0.504) (0.363, 0.653) (0.314, 0.609)
deepseek (0.523, 0.796) (0.146, 0.403) (0.187, 0.457) (0.212, 0.488)
Table 5: Statistical certification bounds on HarmBench illegal (30 scenarios), Claude, gpt-oss, and Llama. Entries are the median of 95% confidence intervals across all specifications under a distribution. Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among the four distributions for each model.
Model RNwJ GPv GPh AwR
claude (0.000, 0.089) (0.005, 0.137) (0.001, 0.106) (0.005, 0.137)
gpt-oss (0.001, 0.106) (0.009, 0.151) (0.013, 0.165) (0.005, 0.137)
llama (0.309, 0.597) (0.086, 0.314) (0.131, 0.382) (0.101, 0.337)
Appendix D ablation study
In this section, we analyze the effect of hyperparameters on certification results. Table  6 shows the hyperparameters and their values used in the experiments. We conduct ablation studies on a randomly selected scenario from the dataset on Graph Path (harmful target constraint) distribution. For Appendices D.5D.7, we certify Llama-3.3-70B-Instruct as they are model-agnostic; otherwise, we certify all evaluated LLMs.
@@ Appendix C Additional Results on HarmBench illegal category
D.1 Jailbreak Probability
Certification bounds on Random Node with Jailbreak distribution is controlled by the jailbreak probability hyperparameter. We show results in Figure 6. Overall, we observe that increasing the jailbreak probability generally raises the certified catastrophic-risk bounds for less robust models such as DeepSeek, Mistral and Llama, with the largest bounds typically appearing at moderate-to-high probabilities (e.g., $p\approx 0.6$$1.0$). For Llama-3.3-70B-Instruct, the bounds increase from $p=0$ to moderate probabilities and then slightly decrease as $p$ approaches $1$, suggesting that overly frequent, highly conspicuous jailbreaks can be partially mitigated by the model. For Claude and gpt-oss, the certified bounds remain relatively low and flat across all probabilities, indicating that these models are comparatively more defensive to the jailbreak prompts used in our experiments.
@@ D.1 Jailbreak Probability
Figure 6: Ablation studies for jailbreak probability on certification bounds for LLMs.
D.2 Length of Query Sequence
Figure 7 shows how certification bounds vary with the length of query sequence. Across models, increasing the sequence length generally pushes the certified bounds upward, indicating that longer conversations provide attackers with more opportunities to elicit catastrophic behavior (for LLaMA, DeepSeek, and Claude). In contrast, gpt-oss appears more robust to query length, with bounds changing only slightly, and for Mistral the bounds also vary little because the model is already highly unsafe even for short sequences.
((a))
((b))
((c))
((d))
((e))
Figure 7: Ablation studies for the length of the query sequence on certification bounds for LLMs.
D.3 Size of Query Set
Figure 8 shows how the size of the query set used to build specifications affects the certified bounds. Across models, increasing the query-set size from 50 to 150 has only a modest effect on the certified bounds. This suggests that, once the initial query set is reasonably large and diverse, our certification results are fairly stable and do not rely on a very specific query-set size.
((a))
((b))
((c))
((d))
((e))
Figure 8: Ablation studies for query-set size on certification bounds for five LLMs.
D.4 Ratio of Weight
To study the influence of this ratio, we perform an ablation experiment by varying $\lambda_{h}/\lambda_{l}$ across the values $\{1.5,2.0,2.5,3.0,3.5\}$. Note that we require $\lambda_{h}>\lambda_{l}$, hence the minimum ratio considered is $1.5$. We then evaluate the resulting certified bounds under these different settings. Figure 9 shows that, for all five LLMs, the certified bounds change only moderately as $\lambda_{h}/\lambda_{l}$ varies, with the highest bounds typically occurring at intermediate ratios (e.g., 2.02.03.03.0). This suggests that a balanced setting—strong enough to move toward the harmful target when queries are accepted, but still willing to step back toward safer neighbors when rejections occur—gives the most effective behavior within this family.
((a))
((b))
((c))
((d))
((e))
Figure 9: Ablation studies for ratio of weights on certification bounds for LLMs.
D.5 Number of Samples
To assess how our certification bounds change with the number of samples $n$, we report bounds in Figure 10(c). The ranges between lower and upper bounds shrink as $n$ increases from small values, and stablize once $n\approx 50$. In our main experiments, we therefore adopt $n=50$ as a trade off between computational cost and statistical precision.
D.6 Graph Thresholds
D.7 Variance
We show the variance of our certification bounds in Figure 10(d), where we run the same experiment on one specification 10 times. We report the median and interquartile range (IQR) of the resulting 95% confidence lower and upper bounds. The results demonstrate that the variance is low, demonstrating the reliability of our certification procedure.
((a))
((b))
((c))
((d))
Figure 10: Ablation studies for (a–b) graph-threshold settings, (c) number of samples, and (d) variance of certified bounds.
Appendix E LLM Usage